# Annabelle Michael Carrell    Albert Gong    Abhishek Shetty    Raaz Dwivedi    Lester Mackey

### v5

@@ Appendix H KT-Compress$(\delta)$
       $\mathcal{X}_{\textup{out}}\leftarrow\mathcal{X}_{\textup{out}}\,\backslash\,\{\boldsymbol{x}\}\cup\{\mathop{\mathrm{missing}}{argmin}_{\boldsymbol{x}^{\prime}\in\mathcal{X}_{\textup{in}}}\operatorname{MMD}_{\mathbf{k}}(\mathbb{P}_{\textup{in}},\mathbb{P}_{\textup{out}}+\frac{1}{n_{\textup{out}}}(\boldsymbol{\delta}_{\boldsymbol{x}^{\prime}}-\boldsymbol{\delta}_{\boldsymbol{x}}))$}
@@
Thm. 4 follows from the following more detailed statement, proved in Sec. I.1 as
@@
Substituting $\mathfrak{M}_{\widetilde{\mathbf{K}}}(\delta,2^{\mathfrak{g}+1}\sqrt{n}_{\textup{in}})=(2^{\mathfrak{g}}\sqrt{n}_{\textup{in}})v_{2^{\mathfrak{g}+1}\sqrt{n}_{\textup{in}},n_{\textup{in}},\widetilde{\mathbf{K}}}\mathopen{}\mathclose{{}\left[\log(\frac{12n_{\textup{in}}4^{\mathfrak{g}}(\log_{4}n_{\textup{in}}-\mathfrak{g})}{2^{\mathfrak{g}+1}\sqrt{n}_{\textup{in}}\delta})}\right]^{-\frac{1}{2}}$ and $C_{\widetilde{\mathbf{K}}}(\delta,2^{\mathfrak{g}+1}\sqrt{n}_{\textup{in}})=(2^{\mathfrak{g}}\sqrt{n}_{\textup{in}})a_{2^{\mathfrak{g}+1}\sqrt{n}_{\textup{in}},n_{\textup{in}},\widetilde{\mathbf{K}}}$ into 223, we obtain the sufficient condition
@@
Suppose Alg is a $(\mathbf{K},\nu,\delta)$-sub-Gaussian thinning algorithm, satisfying Def. 3 with an event $\mathcal{E}$ of probability at least $21-\delta/$ 2. Then Alg is $\mathbf{k}$-sub-Gaussian (Def. I.1) on $\mathcal{E}$ with shift $a_{n_{\textup{out}},n_{\textup{in}},\mathbf{K}}$ and parameter $v_{n_{\textup{out}},n_{\textup{in}},\mathbf{K}}$ defined as
@@
By Props. B.4 and A.1, KH$(\frac{\ell^{2}}{n_{\textup{in}}4^{\mathfrak{g}+1}(\log_{4}n_{\textup{in}}-\mathfrak{g})}\delta)$ with input size $\ell$ and output size $\ell/2$ is a $(\mathbf{K},\nu,\frac{\ell^{2}}{n_{\textup{in}}4^{\mathfrak{g}+1}(\log_{4}n_{\textup{in}}-\mathfrak{g})}\delta)$-sub-Gaussian thinning algorithm with
$\nu\leq\frac{2}{(\ell/2)\sqrt{3}}\sqrt{\log\mathopen{}\mathclose{{}\left(\frac{6(\ell/2)\log_{2}(\ell/(\ell/2))}{\delta}\cdot\frac{n_{\textup{in}}4^{\mathfrak{g}+1}(\log_{4}n_{\textup{in}}-\mathfrak{g})}{\ell^{2}}}\right)\|{\mathbf{K}}\|_{\max}}=\frac{2}{(\ell/2)\sqrt{3}}\sqrt{\log\mathopen{}\mathclose{{}\left(\frac{12n_{\textup{in}}4^{\mathfrak{g}}(\log_{4}n_{\textup{in}}-\mathfrak{g})}{\ell\delta}}\right)\|{\mathbf{K}}\|_{\max}}.$ (228)
@@
Moreover, by the union bound, as detailed in Shetty et al. (2022, App. F.1), every Halve call made by KT-Compress is simultaneously $\mathbf{k}$-sub-Gaussian with these input-size-dependent parameters on a common event of probability at least $1-\frac{\delta}{2}$. Substituting 230 and 231 with $\ell=2^{\mathfrak{g}+1}\sqrt{n}_{\textup{in}}$ into 226, we obtain our error inflation factor expression 218, completing the proof.
@@
Since the deep kernel 43 takes the form
@@
Our final step is to bound the quantile of the sole remaining data-dependent term, $R^{\prime}$. Since the inputs are $c$-sub-Gaussian 44, Lem. 1 of Dwivedi & Mackey (2024) with $\psi^{-1}(r)=\frac{\sqrt{\log r}}{\sqrt{c}}$ implies that the $1-\frac{\widetilde{\beta}}{20s_{n}}$ quantile of $R^{\prime}$ is $O\big{(}\sqrt{\log(\frac{n}{\widetilde{\beta}})}\big{)}$, yielding the result.
@@
Our reasoning is identical to that in App. J with the manifold Gaussian kernel matrix eigenvalue bound 20 now substituted for the Euclidean ball bound 17 and the approximate rank setting $r^{\star}=(\log(nn_{\textup{out}})/c)^{5d^{\star}/2}$ substituted for 240.
@@
L.1 Approximating attention experiment
The experiment of Sec. 4.2 was carried out using Python 3.12.9, PyTorch 2.8.0.dev20250407+cu128 (Paszke et al., 2019), and an Ubuntu 22.04.5 LTS server with an AMD EPYC 7V13 64-Core Processor, 220 GB RAM, and a single NVIDIA A100 GPU (80 GB memory, CUDA 12.8, driver version 570.124.04). For reference, attention layer 1 has $(n,d)=(3136,64)$ and attention layer 2 has $(n,d)=(784,64)$. For each layer and each of the first $50$ ImageNet 2012 validation set batches of size $64$, we measured the time required to complete a forward pass through the layer using CUDA events following $10$ warm-up batches to initialize the GPU. Tab. L.1 provides the hyperparameter settings for each attention approximation in Tab. 3. The settings and implementations for all methods other than Thinformer were provided by Zandieh et al. (2023), and our experiment code builds on their open-source repository https://github.com/majid-daliri/kdeformer.
@@
L.2 Faster SGD training experiment
@@
Technically, the CD-GraB: SBW algorithm requires an a priori upper bound on the maximum Euclidean norm $b_{\max}$ of any stochastic gradient that it will encounter. To conduct our experiment, we first estimate $b_{\max}$ by calculating the maximum gradient Euclidean encountered across $10$ epochs of running SGD with LKH$(\frac{1}{2K})$ reordering. One would typically not choose to carry out such a two-step procedure in practice, but the experiment serves to demonstrate that the CD-GraB: SBW leads to overly conservative performance even if reasonable upper bound is known in advance.
The settings and implementation for both random reshuffling (RR) and CD-GraB: Greedy were those used in the original logistic regression on mortgage application experiment of Cooper et al. (2023). Our experiment code builds on the open-source CD-GraB repository https://github.com/GarlGuo/CD-GraB. As in Cooper et al. (2023), optimization was carried out with a learning rate of $\alpha=0.01$, datapoints were loaded in batches of size $16$, and stochastic gradients were reordered for each datapoint individually.

### v7

@@ Appendix H KT-Compress$(\delta)$
  $\mathcal{X}_{\textup{out}}\leftarrow\mathcal{X}_{\textup{out}}\,\backslash\,\{\boldsymbol{x}\}\cup\{\mathop{\mathrm{argmin}}_{\boldsymbol{x}^{\prime}\in\mathcal{X}_{\textup{in}}}\operatorname{MMD}_{\mathbf{k}}(\mathbb{P}_{\textup{in}},\mathbb{P}_{\textup{out}}+\frac{1}{n_{\textup{out}}}(\boldsymbol{\delta}_{\boldsymbol{x}^{\prime}}-\boldsymbol{\delta}_{\boldsymbol{x}}))$}
@@
Thm. 4 follows from the following more detailed statement, proved in Sec. I.1, as
@@
Substituting $\mathfrak{M}_{\widetilde{\mathbf{K}}}(\delta,2^{\mathfrak{g}+1}\sqrt{n}_{\textup{in}})=(2^{\mathfrak{g}}\sqrt{n}_{\textup{in}})v_{2^{\mathfrak{g}+1}\sqrt{n}_{\textup{in}},n_{\textup{in}},\widetilde{\mathbf{K}}}\mathopen{}\mathclose{{\left[\log(\frac{12n_{\textup{in}}4^{\mathfrak{g}}(\log_{4}n_{\textup{in}}-\mathfrak{g})}{2^{\mathfrak{g}+1}\sqrt{n}_{\textup{in}}\delta})}}\right]^{-\frac{1}{2}}$ and $C_{\widetilde{\mathbf{K}}}(\delta,2^{\mathfrak{g}+1}\sqrt{n}_{\textup{in}})=(2^{\mathfrak{g}}\sqrt{n}_{\textup{in}})a_{2^{\mathfrak{g}+1}\sqrt{n}_{\textup{in}},n_{\textup{in}},\widetilde{\mathbf{K}}}$ into 220, we obtain the sufficient condition
@@
Suppose Alg is a $(\mathbf{K},\nu,\delta)$-sub-Gaussian thinning algorithm, satisfying Def. 3 with an event $\mathcal{E}$ of probability at least $1-\delta/2$. Then Alg is $\mathbf{k}$-sub-Gaussian (Def. I.1) on $\mathcal{E}$ with shift $a_{n_{\textup{out}},n_{\textup{in}},\mathbf{K}}$ and parameter $v_{n_{\textup{out}},n_{\textup{in}},\mathbf{K}}$ defined as
@@
By Props. B.4 and A.1, KH$(\frac{\ell^{2}}{n_{\textup{in}}4^{\mathfrak{g}+1}(\log_{4}n_{\textup{in}}-\mathfrak{g})}\delta)$ with input size $\ell$ and output size $\ell/2$ is a $(\widetilde{\mathbf{K}},\nu,\frac{\ell^{2}}{n_{\textup{in}}4^{\mathfrak{g}+1}(\log_{4}n_{\textup{in}}-\mathfrak{g})}\delta)$-sub-Gaussian thinning algorithm with
$\nu\leq\frac{2}{(\ell/2)\sqrt{3}}\sqrt{\log\mathopen{}\mathclose{{\left(\frac{6(\ell/2)\log_{2}(\ell/(\ell/2))}{\delta}\cdot\frac{n_{\textup{in}}4^{\mathfrak{g}+1}(\log_{4}n_{\textup{in}}-\mathfrak{g})}{\ell^{2}}}}\right)\|{\widetilde{\mathbf{K}}}\|_{\max}}=\frac{2}{(\ell/2)\sqrt{3}}\sqrt{\log\mathopen{}\mathclose{{\left(\frac{12n_{\textup{in}}4^{\mathfrak{g}}(\log_{4}n_{\textup{in}}-\mathfrak{g})}{\ell\delta}}}\right)\|{\widetilde{\mathbf{K}}}\|_{\max}}.$ (225)
@@
Moreover, by the union bound, as detailed in Shetty et al. (2022, App. F.1), every Halve call made by KT-Compress is simultaneously $\mathbf{k}$-sub-Gaussian with these input-size-dependent parameters on a common event of probability at least $1-\frac{\delta}{2}$. Substituting 227 and 228 with $\ell=2^{\mathfrak{g}+1}\sqrt{n}_{\textup{in}}$ into 223, we obtain our error inflation factor expression 215, completing the proof.
@@
Since the deep kernel 40 takes the form
@@
Our final step is to bound the quantile of the sole remaining data-dependent term, $R^{\prime}$. Since the inputs are $c$-sub-Gaussian 41, Lem. 1 of Dwivedi & Mackey (2024) with $\psi^{-1}(r)=\frac{\sqrt{\log r}}{\sqrt{c}}$ implies that the $1-\frac{\widetilde{\beta}}{20s_{n}}$ quantile of $R^{\prime}$ is $O\big{(}\sqrt{\log(\frac{n}{\widetilde{\beta}})}\big{)}$, yielding the result.
@@
Our reasoning is identical to that in App. J with the manifold Gaussian kernel matrix eigenvalue bound 20 now substituted for the Euclidean ball bound 17 and the approximate rank setting $r^{\star}=(\log(nn_{\textup{out}})/c)^{5d^{\star}/2}$ substituted for 237.
@@
L.1 Approximating attention experiments
[…]
The experiment of Tab. 4 was carried out using Python 3.12.9, PyTorch 2.6.0, and an Ubuntu 22.04.5 LTS server with an Intel(R) Xeon(R) Gold 5218 CPU Processor, 100 GB RAM, and a single NVIDIA A6000 GPU (48 GB memory, CUDA 12.1, driver version 530.30.02). For reference, the BigGAN model contains a single attention layer with 409640964096 queries in $\mathbb{R}^{64}$, 102410241024 keys in $\mathbb{R}^{64}$, and 102410241024 values in $\mathbb{R}^{256}$. We measured the time required to complete a forward pass through the approximate softmax matrix 22 layer using CUDA events following 101010 warm-up batches to initialize the GPU. Tab. L.2 provides the hyperparameter settings for each attention approximation in Tab. 4.
The settings and implementations for all methods other than Thinformer were provided by Zandieh et al. (2023), and our experiment code builds on their open-source repository https://github.com/majid-daliri/kdeformer.
@@
Attention Algorithm Layer Configuration
Performer num_features=128
Reformer bucket_size=64
n_hashes=8
ScatterBrain local_context=32
num_features=128
KDEformer sample_size=128
bucket_size=64
Thinformer (Ours) g=2
L.2 Faster SGD training experiments
@@
Technically, the CD-GraB: SBW algorithm requires an a priori upper bound on the maximum Euclidean norm $b_{\max}$ of any stochastic gradient that it will encounter. To conduct our experiment, we first estimate $b_{\max}$ by calculating the maximum gradient Euclidean encountered across 101010 epochs of running SGD with LKH$(\frac{1}{2K})$ reordering. One would typically not choose to carry out such a two-step procedure in practice, but the experiment serves to demonstrate that the CD-GraB: SBW leads to overly conservative performance even if reasonable upper bound is known in advance.
The settings and implementation for both random reshuffling (RR) and CD-GraB: Greedy were those used in the original logistic regression on mortgage application experiment of Cooper et al. (2023). Our experiment code builds on the open-source CD-GraB repository https://github.com/GarlGuo/CD-GraB. As in Cooper et al. (2023), optimization was carried out with a learning rate of $\alpha=0.01$, datapoints were loaded in batches of size 161616, and stochastic gradients were reordered for each datapoint individually.
Figure L.1: Approximate low-rank structure of stochastic gradient matrices. Left. For each epoch $k$ of the LKH$(\frac{1}{2K})$ experiment of Sec. 5.2, we record the $\epsilon_{k}$-rank of the stochastic gradient matrix $\mathbf{X}^{k}\triangleq[\boldsymbol{x}_{1}^{k},\dots,\boldsymbol{x}_{n}^{k}]^{\top}\in\mathbb{R}^{n\times d}$ (see Defs. 4 and 3). Notably, the $\epsilon_{k}$-ranks are significantly smaller than the ambient dimension $d=19$. Right. We display the singular values of the first stochastic gradient matrix, $\mathbf{X}^{1}$. The singular values drop off steeply, resulting in relatively small $\epsilon_{k}$-ranks.
@@
Figure L.2: Slow-growing error inflation. In the experimental setting of Sec. 6.2, the error inflation factor $\widehat{\mathbsf{R}}_{\mathbf{k}_{\textup{deep}}}$ enjoys $O(\log^{5}(n))$ growth as the ratio $\widehat{\mathbsf{R}}_{\mathbf{k}_{\textup{deep}}}/\log^{5}(n)$ decreases with $n$. Here, $\widehat{\mathbsf{R}}_{\mathbf{k}_{\textup{deep}}}^{2}$ is defined by 38 with $n=m$, $n_{\textup{in}}\triangleq\frac{2n}{s}$ and $n_{\textup{out}}\triangleq 2^{\mathfrak{g}}\sqrt{n_{\textup{in}}}$ for constants $\beta=0.05$, $s=32$, and $\mathfrak{g}=4$.